OpenStudy (anonymous):

5x+3y=8 3x+y=8 solve by substitustion

OpenStudy (anonymous):

x = 16

OpenStudy (texaschic101):

3x + y = 8 y = -3x + 8 now sub this into the other equation 5x + 3y = 8 5x + 3(-3x + 8) = 8 5x - 9x + 24 = 8 -- subtract 24 from both sides 5x - 9x = 8 - 24 -- combine like terms -4x = - 16 x = 4 now sub 4 in for x in either equation to find y y = -3x + 8 y = -3(4) + 8 y = -12 + 8 y = -4 check.. 3x + y = 8 3(4) + (-4) = 8 12 - 4 = 8 8 = 8 (correct) x = 4 and y = -4

OpenStudy (anonymous):

thanks again it helped alot